# latexText page

### v1

@@ 4.2.2 Point Game and Segmentation Test
4.2.3 Processing time comparison
In Tab. V, we show the average processing time per image, which counts the total duration from inputting the image and text into CLIP to obtaining the explanation map. Since the gradient can be easily and quickly obtained through the autograd function of Pytorch, both our method and Grad-CAM takes similar processing time as the raw attention and MaskCLIP, which obtain their heat maps from the forward pass of the model and some other minor operations. Note that for gradient-based methods, the backpropagation does not need to go all the way to the input layer, but stops at an intermediate upper layer, and thus the extra computation required is not much. RISE needs the longest processing time, which is a common drawback of perturbation-based methods.
4.3 Ablation study
In this section, we conduct ablation studies to illustrate the influence of the proposed loosened spatial weight (§4.3.1), the number of involved layers (§4.3.2) and multi attention heads (§4.3.3) in the calculation of Grad-ECLIP.
4.3.1 Effect of the loosened spatial weight
Here we conduct ablation study on the $\lambda_{i}$ in Eq. 20 to show the effect of the proposed spatial weight. We consider two versions of Grad-ECLIP with modified spatial importance $\lambda_{i}$: the first version uses the softmax attention rather than 0-1 normalization (denoted as $\lambda_{i}=\mathrm{softmax}$); the second version removes the spatial importance completely by setting $\lambda_{i}=1$.
A comparison of visualizations is presented in Fig. 9. Compared with the heat maps generated by the full Grad-ECLIP in Fig. 9c, the version that removes spatial importance altogether ($\lambda_{i}=1$) contains more noise near object boundaries and on the background (Fig. 9b), but are otherwise consistent with full Grad-ECLIP. The result of using $\lambda_{i}=\mathrm{softmax}$ (Fig. 9a) is equivalent to raw attention (Fig. 3a) due to the output of the $\mathrm{softmax}$ being extremely sparse. We also provide the quantitative comparisons of Grad-ECLIP using $\lambda_{i}=1$ in the quantitative evaluation oin §4.2.1 and §4.2.2, denoted as “w/o $\lambda_{i}$”.
@@ 4.3.1 Effect of the loosened spatial weight
4.3.2 Effect of number of layers on visual explanation
As introduced in §3.2, the explanation can be aggregated over all the layers in Transformer by recursively processing each layer with Eq. 19. In this section, we conduct the experiments to discuss the influence of using different number of layers to generate heat maps for image and text.
With different layer number $N$, the visualizations with specific texts are shown in Fig. 10, and the corresponding caption explanations are shown in Fig. 11. $N=1$ means the visualization is generated only with the final Transformer layer, while $N=12$ means all the layers are involved. The image explanations become worse when increasing the number of layers involved, since the features in lower layer may introduce more noise to the heat map. Therefore, it is the best to just use the last layer in the calculation of image visual explanation, where this conclusion is consistent with the classical gradient-based CAM methods.
As for the text explanation, there is no obvious difference of the visualization quality, since the highlights are basically focusing on “dog”, “car” and “traffic lights” with some minor variations. Therefore, we perform the Deletion and Insertion experiments as in §4.2 on the text explanation maps based on different number of layers $N$. The results are shown in the following Table VI. The explanation faithfulness has the trend that it first increases with more layers used and then goes down with the lower-layer features involved ($N>8$). Therefore, we aggregate the last eight layers maps for interpreting the text encoder in our experiments.
Figure 10: The image visual explanations generated when aggregating over $N$ layers of the image transformer encoder.
Figure 11: The textual explanations generated when aggregating over $N$ layers of the text transformer encoder.
TABLE VI: The text explanation faithfulness vs. the number of transformer layers aggregated for the explanation. Evaluating on MS COCO image-text retrieval (Karpathy’s split) validation dataset: AUC for Deletion and Insertion curves with reporting image retrieval (IR) and text retrieval (TR) performance.
@@ 4.3.2 Effect of number of layers on visual explanation
$N$ IR TR IR TR
1 0.1118 0.2087 0.1059 0.2196
2 0.1021 0.1826 0.1186 0.2351
4 0.0995 0.1786 0.1242 0.2428
6 0.0989 0.1761 0.1273 0.2490
8 0.0996 0.1770 0.1292 0.2536
10 0.1008 0.1843 0.1288 0.2472
12 0.1095 0.2087 0.1219 0.2364
4.3.3 Effect of multi attention heads on visual explanation
@@
As mentioned in (7) of §3.2, for producing the Grad-ECLIP visual explanation, we set CLIP to perform the forward pass with a single head in the attention layer instead of the original multi-head attention layer. In Fig. 12, we show the visualization of explanation maps when using multi-head attention layers, compared to using a single head. Comparing Fig. 12 (a) and (b), using multi-head attention results in some surrounding context information is also highlighted with the explained object.
@@
We further produce the heat maps for each attention head, using the $q\in\mathbb{R}^{D/12}$, $k\in\mathbb{R}^{D/12}$, $v\in\mathbb{R}^{D/12}$, and attention output $o_{cls}\in\mathbb{R}^{D/12}$, where $D$ is channel number before going into multi heads, and visualize them in Fig. 12 (c) for the target “dog”, and (d) for the “car”. The visual explanation in each head highlights different regions, not only seeing the target object. We can infer that the channels assigned to each heads can preserve different information, and the $\mathrm{softmax}$ inside each head helps the model to encode more context information. In contrast, with the single head setting, the $\mathrm{softmax}$ is performed over all channels, which selects out the most important information, and our explanation method can show the model’s attention on the specific explained target, as shown in Fig. 12 (b).
@@
Figure 12: The visual explanation maps with using (a) multi-head attention layer; (b) single-head attention layer; (c) each head in multi-head attention layer for text “dog”; (d) each head in multi-head attention layer for text “car”
5 Analysis of CLIP using Grad-ECLIP
@@
Useful explanation methods can be used to identify failure modes, establish appropriate users’ confidence and give insight to developers to improve models. Therefore, in this section, we use the visual explanation maps and textual explanations generated by Grad-ECLIP to give examples of how to explore the mechanism in text and image matching, and analyze the strengths, weaknesses, and preference of CLIP model. We hope that our explanation tool can help researchers discover more interesting properties of pre-trained image-language models, and inspire further development of these models. Here we analyze three aspects of CLIP: 1) the decomposition and addibility in image-text matching (§5.1); 2) types of attributes that can be identified by CLIP (§5.2); 3) the concreteness/abstractness of words learned by CLIP (§5.3).
@@
5.1 Concept decomposition and addibility in image-text matching
@@
Examining the visualizations shown in Fig. 3h, CLIP can well recognize the single concepts (nouns) and has good attention about actions (verbs). An interesting question is how does it process the combination of words, e.g., adjective and noun, verb and noun? To examine the working function of phrase matching, we conducted experiments comparing the explanation heat maps for single words and combined phrases using Grad-ECLIP.
@@
The results are shown in Fig. 13. Considering adjective-noun combinations in (a), the highlights are put on all three toys when matching with “toy”, and CLIP can successfully highlight the correct toy when the color adjective is included in the text. In the case of “young horse” in (b), the other horses are still highlighted, while the highlights on the young one is strengthened by adding the attribute “young”. The examples of verb-noun cases in (c) and (d) also show similar addibility pattern on the heat maps: (c) with the verb “stand”, the region of person’s leg is highlighted along with the“skis”; (d) with the verb “feed”, the people’s hands are also highlighted together with sheep. We also show some non-existent concepts or strange word combinations in the dashed box of Fig. 13, e.g., “white horse” in (b), “feed skis” in (c), “stand on sheep” in (d). In these cases, the visualization shows that CLIP will mainly focus on the reasonable part of the concept, such as “horse”, “skis” and “sheep”. For the non-existent “white” concept in image (b), the visual explanation does not highlight anything.
@@
Therefore, we infer that when processing the matching of image and phrases, the model has the ability of decomposition and addibility of different concepts. This can help the model to generalize to different scenarios and could be the source of the strong zero-shot ability of CLIP.
@@
Figure 14: Visual explanations on image matching with different kinds of attributions: (a) shape; (b) material; (c) color; (d) size; (e) position. For visualization, the ground-truth corresponding to the text prompt are outlined with white boxes, except for cases involving relative adjectives, e.g.“big”, “small”, “left”, “right”. The text explanation maps are also shown for “small cube” and “left red thing” combinations.
5.2 Diagnostics on attribution identification
@@
In Fig. 13(a), we see that CLIP has an ability to distinguish color attributes, and mark out the corresponding regions on image. To explore further, we conduct an experiment to test CLIP’s ability to identify different types of object attributes. We adopt an example image from CLEVR [79], a diagnostic dataset for visual reasoning, and visualize image-text matching with various attributes: shape (sphere, cylinder, cube), material (metal, matte, plastic), color (red, yellow, blue), size (big, small), position (left, right).
@@
Fig. 14 shows the visual explanation heat maps generated with each image-attribution pair. We have the following findings: 1) for shape and material, the heat maps can show partial correct attention with some obvious objects, such as the metal sphere for “sphere” and the highlighted cylinder and cube for “matte”. However, there are also false positive and false negative errors in (a) and (b). Thus, CLIP possess a certain but limited knowledge about object shapes and materials. 2) For the color attribute in row (c), the results further verify that the model can have good ability to distinguish different colors. 3) For comparative attributes, size (big or small) in (d) and position (left or right) in (e), the visual explanations also show that CLIP produces some erroneous results. For example, there are little differences between the heat maps of “small cube” and “cube” in (a), or “left red thing” and “red” in (c), which demonstrates that the word “cube” and “red” take the major role in the matching. This is also confirmed by the text heat maps in the figure.
@@
Overall, from the above analysis, we infer that CLIP has advantages with common perceptual attributes like color, but cannot well handle physical attributes like shape and material, and is weak at grounding objects with comparative attributes, like size and position relationships. Related to the addibility of concepts in the §5.1, it is reasonable to expect that attributes that have concrete visual appearance, e.g., color, will contribute more to the matching score, compared with the abstract comparative attributes.
Figure 15: The average word importance (via Grad-ECLIP) vs. word concreteness for the top-1000 most frequent words in the MS COCO Karpathy’s validation split. The red dash line shows the linear regression result over the word samples; $r^{2}=0.32,p<0.001$.
@@
5.3 Relationship between word importance and concreteness
From §5.2, we have inferred that the concrete visual concepts (e.g., “red” and “blue”) contribute more to the image and text matching than the abstract attributes (e.g., “left” and “right”). Therefore, based on the text explanation obtained from Grad-ECLIP, we further explore the relationship between word importance in matching and word concreteness, and analyze which type of concepts and words (concrete vs. abstract) that the CLIP model has actually learned and uses most often for matching. For an image-text pair, we calculate the textual explanation from Grad-ECLIP, and then obtain the importance value of each word by normalizing such that the maximum word importance value is $1$ in the sentence. We then calculate the average word importance on the top-1000 most frequent words in the MS COCO caption (Karpathy’s split) validation set. For the word concreteness, we adopt the open sourced database from [80], which provides the concreteness for 40,000 common English words measured by human rating. The concreteness is a value from 1 to 5, where 5 means the most concrete and 1 means the most abstract.
Based on the 1000 selected words, Fig.  15 presents a scatter plot of average word importance versus concreteness value. We perform linear regression analysis on this data (red dashed line), and the regression result was statistically significant ($r^{2}=0.32$, $p<0.001$). The scatter distribution and linear regression result reveal that CLIP places higher word importance to more concrete words, and vice versa, less word importance on more abstract words. Therefore, the words that CLIP has learned for matching are biased towards concrete words.

### v2

@@ IV-B3 User Study on Improving Understanding
IV-B4 Processing time comparison
In Tab. V, we show the average processing time per image, which counts the total duration from inputting the image and text into CLIP to obtaining the explanation map. Since the gradient can be easily and quickly obtained through the autograd function of Pytorch, both our method and Grad-CAM take similar processing time as the raw attention and MaskCLIP, which obtain their heat maps from the forward pass of the model and some other minor operations. Note that for gradient-based methods, the backpropagation does not need to go all the way to the input layer, but stops at an intermediate upper layer, and thus, the extra computation required is not much. RISE needs the longest processing time, which is a common drawback of perturbation-based methods.
IV-C Ablation study
In this section, we conduct ablation studies to illustrate the influence of the proposed loosened spatial weight (§IV-C1), the number of involved layers (§IV-C2), and multi attention heads (§IV-C3) in the calculation of Grad-ECLIP.
IV-C1 Effect of the loosened spatial weight
Here we conduct an ablation study on the $\lambda_{i}$ in Eq. 20 to show the effect of the proposed spatial weight. We consider two versions of Grad-ECLIP with modified spatial importance $\lambda_{i}$: the first version uses the softmax attention rather than 0-1 normalization (denoted as $\lambda_{i}=\mathrm{softmax}$); the second version removes the spatial importance completely by setting $\lambda_{i}=1$.
A comparison of visualizations is presented in Fig. 9. Compared with the heat maps generated by the full Grad-ECLIP in Fig. 9c, the version that removes spatial importance altogether ($\lambda_{i}=1$) contains more noise near object boundaries and on the background (Fig. 9b), but is otherwise consistent with full Grad-ECLIP. The result of using $\lambda_{i}=\mathrm{softmax}$ (Fig. 9a) is equivalent to raw attention (Fig. 3a) due to the output of the $\mathrm{softmax}$ being extremely sparse. We also provide the quantitative comparisons of Grad-ECLIP using $\lambda_{i}=1$ in the §IV-B1 and Appendix C, denoted as “w/o $\lambda_{i}$”.
@@ IV-C1 Effect of the loosened spatial weight
IV-C2 Effect of number of layers on visual explanation
As introduced in §III-B, the explanation can be aggregated over all the layers in Transformer by recursively processing each layer with Eq. 19. In this section, we conduct the experiments to discuss the influence of using different numbers of layers to generate heat maps for images and text.
Let $N_{I}$ and $N_{T}$ be the number of layers used for explaining the image and text encoders, respectively, where $N_{I}=1$ means the visualization was generated with only the final Transformer layer (for image encoder), and $N_{I}=12$ means that the visualizations are aggregated over all Transformer layers. The visualizations for different $N_{I}$ for image explanations are shown in Fig. 10, and the corresponding caption explanations are shown in Fig. 11 for different $N_{T}$. Table VI evaluates the image explanation faithfulness vs. the number of transformer layers ($N_{I}$) aggregated in the image encoder. The setting of $N_{I}=1$ gives the highest IMD. Figure 10 shows the corresponding example for different $N_{I}$ – the lower layers of the image encoder include less semantic information and may introduce more noise to the heat maps, which is why using only the final layer ($N_{I}=1$) in the image encoder has higher faithfulness. Therefore, it is best to just use the last layer in the calculation of image visual explanation, and this conclusion is consistent with the classical gradient-based CAM methods for CNNs.
As for the text explanation, there is no obvious difference in the visualization quality, since the highlights are basically focusing on “dog”, “car”, and “traffic lights” with some minor variations. The faithfulness evaluation results on the text explanation maps based on different numbers of layers $N_{T}$ are shown in the following Table VII. The explanation faithfulness has the trend that it first increases with more layers used and then goes down with the lower-layer features involved ($N_{T}>8$). Therefore, we aggregate the last eight layers of maps for interpreting the text encoder in our experiments.
Figure 10: The image visual explanations generated when aggregating over $N_{I}$ layers of the image transformer encoder.
TABLE VI: The image explanation faithfulness vs. the number of transformer layers aggregated for the explanation. Evaluating on MS COCO image-text retrieval (Karpathy’s split) validation dataset: Deletion, Insertion, and IMD with reporting image retrieval (IR) and text retrieval (TR) performance.
@@ IV-C2 Effect of number of layers on visual explanation
$N_{I}$ IR TR IR TR IR TR
1 0.1246 0.1551 0.1576 0.2056 0.0330 0.0505
2 0.1279 0.1657 0.1547 0.2010 0.0269 0.0253
4 0.1622 0.2051 0.1182 0.1399 -0.0439 -0.0652
8 0.2316 0.2879 0.0484 0.0440 -0.1832 -0.2439
10 0.2121 0.2643 0.0715 0.0774 -0.1407 -0.1868
12 0.2163 0.2642 0.0699 0.0773 -0.1465 -0.1869
Figure 11: The textual explanations generated when aggregating over $N_{T}$ layers of the text transformer encoder.
TABLE VII: The text explanation faithfulness vs. the number of transformer layers aggregated for the explanation. Evaluating on MS COCO image-text retrieval (Karpathy’s split) validation dataset: Deletion, Insertion and IMD with reporting image retrieval (IR) and text retrieval (TR) performance.
[…]
$N_{T}$ IR TR IR TR IR TR
1 0.1118 0.2087 0.1059 0.2196 -0.0059 0.0108
2 0.1021 0.1826 0.1186 0.2351 0.0166 0.0525
4 0.0995 0.1786 0.1242 0.2428 0.0247 0.0643
6 0.0989 0.1761 0.1273 0.2490 0.0284 0.0729
8 0.0996 0.1770 0.1292 0.2536 0.0296 0.0766
10 0.1008 0.1843 0.1288 0.2472 0.0279 0.0629
12 0.1095 0.2087 0.1219 0.2364 0.0125 0.0277
IV-C3 Effect of multi attention heads on visual explanation
@@
As mentioned in (7) of §III-B, for producing the Grad-ECLIP visual explanation, we set CLIP to perform the forward pass with a single head in the attention layer instead of the original multi-head attention layer. In Fig. 12, we show the visualization of explanation maps when using multi-head attention layers, compared to using a single head. Comparing Fig. 12 (a) and (b), using multi-head attention results in some surrounding context information being highlighted with the explained object.
@@
We further produce the heat maps for each attention head, using the $q\in\mathbb{R}^{D/12}$, $k\in\mathbb{R}^{D/12}$, $v\in\mathbb{R}^{D/12}$, and attention output $o_{cls}\in\mathbb{R}^{D/12}$, where $D$ is channel number before going into multi heads, and visualize them in Fig. 12 (c) for the target “dog”, and (d) for the “car”. The visual explanation in each head highlights different regions, not only seeing the target object. We can infer that the channels assigned to each head can preserve different information, and the $\mathrm{softmax}$ inside each head helps the model to encode more context information. In contrast, with the single head setting, the $\mathrm{softmax}$ is performed over all channels, which selects out the most important information, and our explanation method can show the model’s attention on the specific explained target, as shown in Fig. 12 (b).
@@
Figure 12: The visual explanation maps using (a) multi-head attention layer; (b) single-head attention layer; (c) each head in the multi-head attention layer for text “dog”; (d) each head in the multi-head attention layer for text “car”.
V Analysis of CLIP using Grad-ECLIP
@@
Useful explanation methods can be used to identify failure modes, establish appropriate users’ confidence, and give insight to developers to improve models. Therefore, in this section, we use the visual explanation maps and textual explanations generated by Grad-ECLIP to give examples of how to explore the mechanism in text and image matching, and analyze the strengths, weaknesses, and preferences of the CLIP model. We hope that our explanation tool can help researchers discover more interesting properties of pre-trained image-language models and inspire further development of these models. Here we analyze three aspects of CLIP: 1) the decomposition and addibility in image-text matching (§V-A); 2) types of attributes that can be identified by CLIP (§V-B); 3) the concreteness/abstractness of words learned by CLIP (§V-C). We note that there is a risk that the heat map explanations may not be accurate when explaining the model, and we can temper this risk by selecting the XAI that has higher faithfulness to the model. We have shown in Section IV that our method Grad-ECLIP has the highest faithfulness among the related methods, and thus, this provides a justification for using it to perform our interpretation/analysis of CLIP here. Future XAI methods may improve on faithfulness and may update and refine these interpretations. Nonetheless, this is the nature of scientific progress.
@@
V-A Concept decomposition and addibility
@@
Examining the visualizations shown in Fig. 3h, CLIP can well recognize the single concepts (nouns) and has good attention to actions (verbs). An interesting question is how it processes the combination of words, e.g., adjective and noun, verb and noun? To examine the working function of phrase matching, we conducted experiments comparing the explanation heat maps for single words and combined phrases using Grad-ECLIP.
@@
The results are shown in Fig. 13. Considering adjective-noun combinations in (a), the highlights are put on all three toys when matching with “toy”, and CLIP can successfully highlight the correct toy when the color adjective is included in the text. In the case of “young horse” in (b), the other horses are still highlighted, while the highlights on the young one are strengthened by adding the attribute “young”. The examples of verb-noun cases in (c) and (d) also show a similar addibility pattern on the heat maps: (c) with the verb “stand”, the region of the person’s leg is highlighted along with the“skis”; (d) with the verb “feed”, the people’s hands are also highlighted together with sheep. We also show some non-existent concepts or strange word combinations in the dashed box of Fig. 13, e.g., “white horse” in (b), “feed skis” in (c), “stand on sheep” in (d). In these cases, the visualization shows that CLIP will mainly focus on the reasonable part of the concept, such as “horse”, “skis” and “sheep”. For the non-existent “white” concept in image (b), the visual explanation does not highlight anything.
To verify the above observation quantitatively, we conduct an experiment on 60 randomly selected images from the MS COCO validation set. Concretely, for each image, we provide three words based on the content of the image, including one base word, one mismatched word, and one matched word. For example, in Figure. 13(b), the base word is “horse”, and the matched and mismatched words are “young” and “white”, respectively. Combining the base word with the matched word or mismatched word produces the matched phrase, e.g. “young horse”, and the mismatched phrase “white horse”, respectively. We randomly select 30 images for the combination of adjective and noun, and another 30 images for verb and noun combinations.
To measure concept addibility, we calculate the cosine similarity between the explanation heat map for the whole phrase, and the combination of corresponding word-based heat maps, e.g., the heat map for the matched phrase “young horse” and the combination of maps for “young” and for “horse”. The averaged similarities over all images are shown in Tab. VIII , where we list the results using different combination methods. Combining word-based maps using addition leads to the highest similarity with the phrase-based heat map (p<<.001, paired t-test between addition and other methods), regardless of whether it is a matched or mismatched phrase. Meanwhile, combinations using multiplication have significantly lower similarity compared to addition, which supports the above addibility hypothesis observed from Fig. 13.
TABLE VIII: Average cosine similarity between the heat map of a phrase and the combined heat maps of individual words using different combination methods. The last column shows the p-value for a paired sample t-test between “add maps” and other methods.
Matched Phrase Mismatched Phrase p-value
Combination method adj.+noun verb+noun adj.+noun verb+noun vs. add maps
add maps 0.9750 0.9814 0.9480 0.9755 -
multiply maps 0.7871 0.7231 0.5810 0.5400 <<.001
maximum maps 0.9619 0.9741 0.9319 0.9711 <<.001
Figure 14: (Left) Scatter plot of explanation heat map similarity vs. CLIP image-text matching score for matched words, mismatched words, matched phrases, and mismatched phrases. The heat map similarity is with the base word, and the CLIP score is relative to the CLIP score using the base word. Thus, all base word points are at coordinates (1,1). (Right) Scatter plot and linear regression of the relative CLIP image-phrase matching score vs. the relative image-word matching score for mismatched words (top-right) and matched words (bottom-right).
We conducted a further analysis by plotting the heat map similarity of the matched/mismatched word or phrase compared to the base word, versus the CLIP image-text matching score, relative to the base word, in Fig. 14(left). The mismatched phrase (red “+”) usually obtains a lower CLIP matching score than the base word (the yellow star) but a higher score than the mismatched word (purple circle), while the explanation heat maps of mismatched phrases have high similarity with the base word (near 1). Therefore, combining with the verified addibility pattern, we infer that when CLIP matches the image with a phrase, it separates the phrase into words and makes the matching result based on putting attention to each word. Thus, for the case of a mismatched phrase (e.g., the white horse in Fig. 13(b)) where the object does not exist in the image, the model basically looks at the image region corresponding to the base word “horse” and makes the prediction. The scatter plot in Fig. 14 (top-right) also verifies that the CLIP matching score for mismatched words is usually lower than the base words (lower than 1), while the scores for the corresponding mismatched phrases are close to 1, i.e., similar to just base-word matching score. While there is a linear trend between the relative image-mismatched-word score and the image-phrase score ($R^{2}=0.11$, $p=0.007$ in linear regression), the $R^{2}$ value is extremely low, indicating significant unexplained variance. On the other hand, in Fig. 14 (bottom-right), most matched phrases will have increased score compared to the baseline word, and there is a strong linear trend ($R^{2}=0.61$, $p<.001$), where matched words with higher scores will also have higher scores in the matched phrase. A matched word can provide further positive effect to using just the base word in the matching process, so that most matched phrases (green “+”) obtain higher CLIP scores than the base word (the yellow star).
@@
Therefore, we infer that when processing the matching of the image and phrases, the model has the ability of decomposition and addibility of different concepts. This can help the model to generalize to different scenarios and could be the source of the strong zero-shot ability of CLIP.
@@
Figure 15: Visual explanations on image matching with different kinds of attributions: (a) shape; (b) material; (c) color; (d) size; (e) position. For visualization, the ground-truth corresponding to the text prompt is outlined with white boxes, except for cases involving relative adjectives, e.g.“big”, “small”, “left”, “right”. The text explanation maps are also shown for “small cube” and “left red thing” combinations.
V-B Diagnostics on attribution identification
@@
In Fig. 13(a), we see that CLIP has an ability to distinguish color attributes and mark out the corresponding regions on the image. To explore further, we conduct an experiment to test CLIP’s ability to identify different types of object attributes. We adopt an example image from CLEVR [22], a diagnostic dataset for visual reasoning, and visualize image-text matching with various attributes: shape (sphere, cylinder, cube), material (metal, matte, plastic), color (red, yellow, blue), size (big, small), position (left, right).
@@
Fig. 15 shows the visual explanation heat maps generated with each image-attribution pair. We have the following findings: 1) for shape and material, the heat maps can show partial correct attention with some obvious objects, such as the metal sphere for “sphere” and the highlighted cylinder and cube for “matte”. However, there are also false positive and false negative errors in (a) and (b). Thus, CLIP possesses a certain but limited knowledge about object shapes and materials. 2) For the color attribute in row (c), the results further verify that the model can have a good ability to distinguish different colors. 3) For comparative attributes, size (big or small) in (d) and position (left or right) in (e), the visual explanations also show that CLIP produces some erroneous results. For example, there are a few differences between the heat maps of “small cube” and “cube” in (a), or “left red thing” and “red” in (c), which demonstrates that the words “cube” and “red” take the major role in the matching. This is also confirmed by the text heat maps in the figure.
Figure 16: The ground truth objects (outlined in white) and the Grad-ECLIP explanation heat maps corresponding to image-text match for specific attribute words: (a) Color “red object”; (b) Shape “sphere object”; (c) Material “rubber object”; (d) Size “small object”; (e) Location relationship “left object”.
TABLE IX: Averaged similarity/distance between the ground-truth mask and Grad-ECLIP heat maps generated when CLIP matches attribute words in different categories.
Color Shape Material Size Location
Cosine similarity $\uparrow$ 0.6654 0.5959 0.5772 0.4645 0.2758
Correlation $\uparrow$ 0.0419 0.0218 0.0150 0.0100 0.0107
OTD $\downarrow$ 0.0085 0.0084 0.0090 0.0093 0.0356
KLD $\downarrow$ 2.0237 2.0968 3.1444 4.3012 5.0392
JSD $\downarrow$ 0.2881 0.3222 0.3072 0.3856 0.5327
We next conduct a quantitative experiment to explore CLIP’s ability to encode different object attributes. We utilize the annotations of the CLEVR dataset, which provides the center position of each object with its properties, including color, shape, material, size, and location relationship. With the object coordinates, we use SAM [24] to generate segmentation masks for the objects, and then perform manual checking to obtain the ground-truth masks. We randomly select 50 images from the dataset. An image example with the ground truth masks and explanation heat maps for specific attribute text is shown in Fig. 16. Using different attribute words, we measure similarity or distance between the CLIP explanation maps and the corresponding ground-truth, with higher similarity (lower distance) indicating that CLIP encodes relevant features for that attribute by focusing on the corresponding object.
The results averaged over attribute categories are presented in Tab. IX. The higher average similarities (or lower distances) for color words demonstrate that the attention of CLIP can better locate the correct object when matching with the given object colors, which means CLIP has a better ability to distinguish or encode color attributes. Following color, the shape and material attributes are the 2nd and 3rd best, while size and location are the worst attributes. To further verify the bad matching with location relationship words, we further conduct an ablation study by comparing using only shape, e.g., “sphere”, the combination of location and shape, e.g., “left sphere”, and only location, “e.g., left”, to describe the left object. The results show that the location relationship words bring no influence to the CLIP matching process, with the heatmap similarity unchanged when adding the location attribute word. See Appendix B for details.
@@
Overall, from the above analysis, we infer that CLIP has advantages with common perceptual attributes like color, but cannot well handle physical attributes like shape and material, and is weak at grounding objects with comparative attributes, like size and position relationships. Related to the addibility of concepts in the §V-A, it is reasonable to expect that attributes that have a concrete visual appearance, e.g., color, will contribute more to the matching score, compared with the abstract comparative attributes.
@@
Figure 18: The average word importance (via Grad-ECLIP) vs. word concreteness for the top-1000 most frequent words in the MS COCO Karpathy’s validation split. The red dashed line shows the linear regression result over the word samples; $r^{2}=0.32,p<0.001$.
V-C Relationship between word importance and concreteness
From §V-B, we have inferred that the concrete visual concepts (e.g., “red” and “blue”) contribute more to the image and text matching than the abstract attributes (e.g., “left” and “right”). Therefore, based on the text explanation obtained from Grad-ECLIP, we further explore the relationship between word importance in matching and word concreteness, and analyze which type of concepts and words (concrete vs. abstract) the CLIP model has actually learned and uses most often for matching. For an image-text pair, we calculate the textual explanation from Grad-ECLIP, and then obtain the importance value of each word by normalizing such that the maximum word importance value is $1$ in the sentence. We then calculate the average word importance on the top-1000 most frequent words in the MS COCO caption (Karpathy’s split) validation set. For the word concreteness, we adopt the open-sourced database from [5], which provides the concreteness for 40,000 common English words measured by human rating. The concreteness is a value from 1 to 5, where 5 means the most concrete and 1 means the most abstract.
Based on the 1000 selected words, Fig.  18 presents a scatter plot of average word importance versus concreteness value. We perform linear regression analysis on this data (red dashed line), and the regression result was statistically significant ($r^{2}=0.32$, $p<0.001$). The scatter distribution and linear regression result reveal that CLIP places higher word importance on more concrete words, and vice versa, less word importance on more abstract words. Therefore, the words that CLIP has learned for matching are biased towards concrete words.